# ADPro: a Test-time Adaptive Diffusion Policy for Robot Manipulation via Manifold and Initial Noise Constraints

### v1

@@ 4.3 Generalization on CALVIN
Table 3: Evaluation of generalizability of ADPro on RealWP for five continuous key pose prediction, with and without fine-tuning.
@@
Table 4: Comparison of grasp success and inference time across 4 Acronym object categories.
@@
4.4 Generalization on Real-world Data
@@
Real_World_Play (RealWP) (Mees, Borja-Diaz, and Burgard 2023) contains 9 hours of unstructured real-world data collected by experts using Franka Panda robots. The dataset covers over 25 distinct manipulation skills, which are highly diverse and challenging due to their unstructured and sometimes suboptimal nature. A total of 3,605 episodes are annotated with language descriptions of the robot’s behavior. Each episode includes RGB-D images from a static and a gripper-mounted camera, proprioceptive data, and a 7-DoF action. Since RealWP lacks a motion planner, we predict gripper pose trajectories, similar to CALVIN. To do this, we extract frames with significant motion changes as key poses using the algorithm from 3D Diffuser Actor. This process yields a trajectory of 6, 7, or 8 steps for each episode.
@@
To evaluate the generalization ability of ADP, we directly applied the model trained on CALVIN to predict five key poses sequentially. The accuracy of consecutive five-step trajectory, shown on the right side of Tab. 3, demonstrates that our method enhances generalization across tasks and environments, even without fine-tuning. We further fine-tuned the model using 80% of the 3,605 episodes and evaluated it on the remaining 20%. The results, presented on the left Tab. 3, show that after fine-tuning, our ADPro substantially boosts the policy’s one-step success rate by 20 percentage points.
4.5 Ablations
@@
The proposed adaptive policy can be seamlessly integrated into existing diffusion policies in a plug-and-play manner. To assess its effectiveness in improving success rate and efficiency of manipulation, we incorporated it into three baseline models: DP3, SE(3)-DiffusionFields (SEDF) (Urain et al. 2023), and Diffuser. We then conducted a quantitative comparison of their results, as shown in Tab. 2 and 4.
@@
For SEDF, we evaluate the pre-trained model on the Acronym dataset (Eppner, Mousavian, and Fox 2021), which provides successful 6-DoF grasps for diverse ShapeNet (Chang et al. 2015) objects. Grasp success rates are evaluated in simulation using Isaac Gym (Makoviychuk et al. 2021), following the same evaluation protocol of (Urain et al. 2023). We evaluate the model on four representative object categories: Bottle, Laptop, Book, and ToyFigure, each with the first 50 instances selected from the dataset. For each category, we report the average grasp success rate and the average generation time per grasp (in seconds), under both low-step (S=20) and full-step (S=120) diffusion settings. The results are summarized in Tab. 4. Our method (SEDF+ADP) achieves either comparable or superior success rates compared to both the baseline SEDF (S=20) and the high-step SEDF (S=120), while achieving up to 5× faster grasp generation compared to SEDF (S=120). In addition, we evaluate the multi-object model on a diverse set of 61 objects, each from a distinct category, by executing 200 generated grasps per object. As shown in Fig. 6, integrating ADP consistently improves performance across all diffusion step settings. On average, our method achieves a 2.7% increase in success rate compared to SEDF with the same steps.
@@
Furthermore, to evaluate the effectiveness of the major component of our ADPro, we consider the following ablative versions of our model: i) without (w/o) initial noise constraint (INC), ii) w/o spherical Gaussian constraint (SGC), iii) w/o observation guidance (OG). Ablation studies on the RLBench and CALVIN benchmarks show that all three modules contribute effectively to model performance. Notably, removing the OG module leads to a substantial drop in performance and increased time costs, as shown in Tab. 5.
@@
Table 5: Ablation study. Our model significantly outperforms its counterparts that do not use the initial noise constraint, the Gaussian manifold constraint, or the observation guidance.

### v2

@@ IV-C Generalization on CALVIN
TABLE III: Evaluation of generalizability of ADPro on RealWP for five continuous key pose prediction, with and without fine-tuning.
@@
TABLE IV: Comparison of grasp success and inference time across 4 Acronym object categories.
@@
IV-D Generalization on Real-world Data
@@
Real_World_Play (RealWP) [87] contains 9 hours of unstructured real-world data collected by experts using Franka Panda robots. The dataset covers over 25 distinct manipulation skills, which are highly diverse and challenging due to their unstructured and sometimes suboptimal nature. A total of 3,605 episodes are annotated with language descriptions of the robot’s behavior. Each episode includes RGB-D images from a static and a gripper-mounted camera, proprioceptive data, and a 7-DoF action. Since RealWP lacks a motion planner, we predict gripper pose trajectories, similar to CALVIN. To do this, we extract frames with significant motion changes as key poses using the algorithm from 3D Diffuser Actor. This process yields a trajectory of 6, 7, or 8 steps for each episode.
@@
To evaluate the generalization ability of ADP, we directly applied the model trained on CALVIN to predict five key poses sequentially. The accuracy of the five consecutive steps trajectory, shown on the right side of Tab. III, demonstrates that our method enhances generalization across tasks and environments, even without fine-tuning. We further fine-tuned the model using 80% of the 3,605 episodes and evaluated it on the remaining 20%. The results, presented on the left Tab. III, show that after fine-tuning, our ADPro substantially boosts the policy’s one-step success rate by 20 percentage points.
IV-E Ablations
@@
The proposed adaptive policy can be seamlessly integrated into existing diffusion policies in a plug-and-play manner. To assess its effectiveness in improving success rate and efficiency of manipulation, we incorporated it into three baseline models: DP3, SE(3)-DiffusionFields (SEDF) [88], and Diffuser. We then conducted a quantitative comparison of their results, as shown in Tab. II and IV.
@@
For SEDF, we evaluate the pre-trained model on the Acronym dataset [89], which provides successful 6-DoF grasps for diverse ShapeNet [90] objects. Grasp success rates are evaluated in simulation using Isaac Gym [91], following the same evaluation protocol of [88]. We evaluate the model on four representative object categories: Bottle, Laptop, Book, and ToyFigure, each with the first 50 instances selected from the dataset. For each category, we report the average grasp success rate and the average generation time per grasp (in seconds), under both low-step (S=20) and full-step (S=120) diffusion settings. The results are summarized in Tab. IV. Our method (SEDF+ADP) achieves either comparable or superior success rates compared to both the baseline SEDF (S=20) and the high-step SEDF (S=120), while achieving up to 5× faster grasp generation compared to SEDF (S=120). In addition, we evaluate the multi-object model on a diverse set of 61 objects, each from a distinct category, by executing 200 generated grasps per object. As shown in Fig. 6, integrating ADP consistently improves performance across all diffusion step settings. On average, our method achieves a 2.7% increase in success rate compared to SEDF with the same steps.
@@
Furthermore, to evaluate the effectiveness of the major component of our ADPro, we consider the following ablative versions of our model: i) without (w/o) initial noise constraint (INC), ii) w/o spherical Gaussian constraint (SGC), iii) w/o observation guidance (OG). Ablation studies on the RLBench and CALVIN benchmarks show that all three modules contribute effectively to model performance. Notably, removing the OG module leads to a substantial drop in performance and increased time costs, as shown in Tab. V.
@@
TABLE V: Ablation study. Our model significantly outperforms its counterparts that do not use the initial noise constraint, the Gaussian manifold constraint, or the observation guidance.